# thanks: This paper has been submitted in conjunction with its companion Algeri et al. [1]

### v1

@@ I Introduction
Estimating and testing models for the angular power spectrum is required in many astrophysical and cosmological settings, including the cosmic microwave background [2], galaxy sky surveys [3], and weak gravitational lensing surveys [4]. The focus here is on the stochastic gravitational-wave background (SGWB) resulting from the superposition of gravitational-wave (GW) signals that are too weak or too numerous to be resolved individually. While the SGWB and its anisotropies have not yet been detected by second-generation GW detectors such as the Advanced LIGO [5] and Advanced Virgo [6], upper limits have been placed on the GW energy density of the background as well as its angular power spectra [7, 8].
@@
As shown in Appendix A, the estimators of the angular power spectrum follow a generalized $\chi^{2}$ distribution, which cannot be accurately approximated by the Gaussian distribution. Moreover, using the generalized $\chi^{2}$ likelihood is nontrivial because, in general, the density does not have a closed-form expression. Alternatively, one could devise a scheme that does not require any assumptions regarding the angular power spectrum’s underlying distribution. The latter approach is used here to implement and adapt the statistical procedure outlined in our companion paper [1] to the context of the SGWB.
@@ II Spherical Harmonics Search Methods for the SGWB
GW data from the Advanced LIGO and Advanced Virgo detectors may be used to generate estimators of the SGWB angular power spectrum. Each GW detector, labeled by $I=1,2$, produces a strain time series given by
$s_{I}(t)=h_{I}(t)+n_{I}(t),$
with $n_{I}(t)$ denoting the detector noise and $h_{I}(t)$ is the detector response to a GW signal. The GW signal can be decomposed in terms of GW modes $h_{A}(f,\hat{\Omega})$ defined by their polarization $A$, frequency $f$, and propagation direction $\hat{\Omega}$:
$h_{I}(t)=\int^{\infty}_{-\infty}\mathrm{d}f\int_{S^{2}}\mathrm{d}\hat{\Omega}h_{A}(f,\hat{\Omega})F_{I}^{A}(\hat{\Omega},t)e^{i2\pi f(t-\hat{\Omega}\cdot\vec{x}_{I}(t)/c)}.$
@@
$F_{I}^{A}(\hat{\Omega},t)=\frac{1}{2}\left[\hat{X}^{a}_{I}(t)\hat{X}^{b}_{I}(t)-\hat{Y}^{a}_{I}(t)\hat{Y}^{b}_{I}(t)\right]e^{A}_{ab}(\hat{\Omega}),$
in which we assume Einstein summation over the spatial indices $a,b$. The gravitational-wave polarization tensors $e^{A}_{ab}(\hat{\Omega})$ are defined as in [23]. We assume the SGWB is unpolarized, Gaussian, and stationary. Additionally, by assuming perturbations to the spacetime metric are real, the Fourier coefficients $h_{A}(f,\hat{\Omega})$ satisfy $h_{A}(-f,\hat{\Omega})=h_{A}^{*}(f,\hat{\Omega})$, where denotes complex conjugation. They also have zero means, $\langle h_{A}(f,\hat{\Omega})\rangle=0$, and the expected value of their two-point correlation is
$\langle h_{A}^{*}(f,\hat{\Omega})\,h_{A^{\prime}}(f^{\prime},\hat{\Omega}^{\prime})\rangle=\frac{1}{4}\,\mathcal{P}(f,\hat{\Omega})\,\delta_{AA^{\prime}}\,\delta(f-f^{\prime})\,\delta^{2}(\hat{\Omega},\hat{\Omega}^{\prime}),\,$
where ${\mathcal{P}}(f,\hat{\Omega})$ gives the spectral and angular distribution of the SGWB [23], $\delta_{AA^{\prime}}$ denotes the Kronecker delta function in polarization, and the remaining $\delta$’s denote Dirac delta functions in frequency and angular position on the sky. We also assume that ${\mathcal{P}}(f,\hat{\Omega})$ can be factored into its separate spectral and angular components:
$\mathcal{P}(f,\hat{\Omega})=H(f)\,\mathcal{P}(\hat{\Omega}),$
@@
$H(f)=\left(\frac{f}{f_{\rm ref}}\right)^{\alpha-3}.$
We set the reference frequency $f_{\rm ref}=25$ Hz while the spectral index $\alpha$ may be assigned a value depending on the type of SGWB model being considered. For the astrophysical SGWB resulting from merger events between compact objects such as black holes and neutron stars, $\alpha=2/3$ [7].
Next, decompose the angular distribution of the background $\mathcal{P}(\hat{\Omega})$ into its spherical harmonic components $Y_{\ell m}$ and coefficients $a_{\ell m}$:
$\mathcal{P}(\hat{\Omega})=\sum_{\ell=0}^{\ell_{\rm max}}\sum_{m=-\ell}^{\ell}a_{\ell m}\,Y_{\ell m}(\hat{\Omega}).$
Note that these definitions assume that each GW frequency bin has the same anisotropy. This assumption can be relaxed, and the analysis can be pursued in separate frequency bins (or bands) ${f_{b}}$. In this case, the spherical harmonic coefficients will become frequency dependent and we denote them as $a_{{f_{b}},\ell m}$.
The value of $\ell_{max}$ determines the highest-order spherical harmonic modes that are considered in the analysis. This value may be chosen differently for different spectral models of the SGWB, and while $\ell_{max}=4$ is typically used when $\alpha=2/3$ [7], $\ell_{max}=8$ is used here. This choice of $\ell_{max}$ is more appropriate for the frequency range over which we will evaluate the angular power spectrum. It has been shown that while $\ell_{max}=4$ is an optimal choice for making a GW detection of the chosen SGWB model, it corresponds to an angular resolution that is lower than what is resolvable by the Advanced LIGO-Virgo detectors [30], which justifies considering spherical harmonic modes corresponding to a higher angular resolution.
@@
[…]
@@
$C(t;f)=\frac{2}{\tau}\tilde{s}_{I}^{*}(t;f)\tilde{s}_{J}(t;f),$
where, to simplify the notation, the $I,J$ indices corresponding to the two detectors from the cross-spectra terms have been dropped. Assuming that the noise of the two detectors is uncorrelated [23], the expectation value of the cross-correlation spectrum can be expressed as
$\langle C(t;f)\rangle=\sum_{\ell=0}^{\ell_{\rm max}}\sum_{m=-\ell}^{\ell}H(f)\gamma_{\ell m}(t;f){a}_{{f_{b}},\ell m},$
@@
$N_{ft,f^{\prime}t^{\prime}}=\langle C(t;f)C^{*}(t^{\prime};f^{\prime})\rangle-\langle C(t;f)\rangle\langle C^{*}(t^{\prime};f^{\prime})\rangle$
@@
Here, $P_{I}(t;f)$ and $P_{J}(t;f)$ correspond to the one-sided power spectral densities of the two GW detectors’ outputs. Assuming the present GW signal powers (both cross- and auto-correlated) are much less than the detector noise power, we obtain the approximation in Eq. (1). For a given set of $a_{{f_{b}},\ell m}$, the likelihood function for the cross-correlation spectra in the frequency band $f_{b}$ is proportional to
$\exp(-\sum_{{f},t}[C(t;{f})-H({f})\,\sum_{\ell m}\gamma_{\ell m}(t;{f})\,{a}_{{f_{b}},\ell m}]^{*}$ (2)
$\quad N^{-1}_{{f}t,{f}t}[C(t;{f})-H({f})\,\sum_{\ell m}\gamma_{\ell m}(t;{f})\,{a}_{{f_{b}},\ell m}])$
@@
Here, $\Gamma$ is the Fisher information matrix [23]. Note that if Eq. (2) holds, each $\hat{a}_{{f_{b}},\ell m}$ is a linear combination of complex Gaussian variables; thus, they also follow a complex Gaussian distribution with mean $\langle\hat{a}_{{f_{b}},\ell m}\rangle={a}_{{f_{b}},\ell m}$ (discussed further below) and covariance $\operatorname{Cov}(\hat{a}_{{f_{b}},\ell m},\hat{a}_{{f_{b}},\ell^{\prime}m^{\prime}})=(\Gamma^{-1}_{{f_{b}}})_{\ell m,\ell^{\prime}m^{\prime}}.$ Indeed the same is true, at least approximately, even if Eq. (2) does not hold, but the summation over $t$ in Eq. (4) is taken over a large number of segments (as is the case with the Advanced LIGO and Advanced Virgo observing runs).
Eq. (4) defines a map of the gravitational-wave background convolved with the antenna pattern of the detector pair, the so-called dirty map. The covariance matrix of the dirty map is given in Eq. (5). Since Eq. (3) represents the deconvolution of the GW signal from the detector response, it is the so-called clean map. The covariance matrix of the clean map is obtained by inverting the covariance matrix of the dirty map.
@@
From a statistical perspective, maximum likelihood is fundamentally a method for estimating deterministic quantities. This means the $\hat{a}_{{f_{b}},\ell m}$ in Eq. (3) serve as maximum likelihood estimators of the spherical harmonic coefficients, $a_{{f_{b}},\ell m}$, that uniquely characterize the SGWB of the only realization of the universe available. Therefore, similarly to what is implicitly done in many statistical analyses, the $a_{{f_{b}},\ell m}$ are not treated as random variables [e.g., 23, 31, 32, 7]. This is a reasonable assumption in practical scenarios, and treating the $a_{{f_{b}},\ell m}$ as fixed avoids the risk of incorrectly specifying their underlying distribution, which can substantially bias the resulting inference.
The proposed statistical methodology requires that multiple replicates of $\hat{a}_{{f_{b}},\ell m}$ are available. To ensure that is the case, we partition the total observation time $T$ into $S$ segments, each of duration $\Delta T$, so that $T=\Delta T\cdot S.$ The third observing run (O3) includes data from both Advanced LIGO detectors as well as the Advanced Virgo detector, collected from April 1, 2019, to March 27, 2020, with a month-long pause in October 2019. Aside from this pause, the duty cycles of the three detectors were 77%, 75%, and 76% for LIGO Livingston, LIGO Hanford, and Virgo, respectively [33]. For this study, we subdivide the O3 data set into $S=$15 segments, each corresponding to approximately 10 days of detector live time. We also divide the frequency range 20-180 Hz into $B=8$ discrete frequency bins, each 20 Hz wide. We assume the same time-domain and frequency-domain data quality cuts that were applied during other O3 anisotropic SGWB search analyses [7].
@@
The derivation of Eq. (7) is given in Appendix B. Previous works [e.g., 26] have omitted the second term in the numerator of Eq. (7), implicitly making the assumption that $a_{{f_{b}},\ell m}$’s are complex Gaussian random variables with zero mean. However, this assumption is not consistent with estimating the coefficients $a_{{f_{b}},\ell m}$ (as realized in our universe) via the maximum likelihood technique – that is, likelihood maximization implicitly assumes the coefficients to be the (deterministic) limit to which the maximum likelihood estimates converge. In this manuscript, this inconsistency is avoided by treating $a_{{f_{b}},\ell m}$ as fixed.
Finally, it is worth mentioning that one could always replace ${\Gamma}_{{f_{b}},s}^{-1}$ in Eq. (7) with its regularized counterpart if a regularization scheme were applied to invert the Fisher matrix [Cf. 23, 26, 35]. Such regularization is typically implemented via singular value decomposition to mitigate small singular values arising from directions in the sky to which the detector network is insensitive [23, 7]. However, this procedure introduces a bias that propagates into $\Sigma_{{f_{b}},s}$ in nontrivial ways [27]. To avoid this bias, the Fisher matrix is not regularized here.
@@ III Statistical Methods
Let $\bm{\theta}\in\mathbb{R}^{p}$ be an unknown parameter vector, and denote with $\bm{A}_{f_{b}}(\bm{\theta})$ the proposed model for the angular power spectrum. Now, assume that the $\hat{\bm{A}}_{{f_{b}},s}$ vectors are independent across all frequency bins and data segments. Hence, in this context, ‘large sample’ asymptotic results refer to the situation in which $N=\ell_{\rm max}\cdot B\cdot S\rightarrow\infty$, but, since typically the $\ell_{\rm max}$ elements of the vectors $\hat{\bm{A}}_{{f_{b}},s}$ are not independent for different values of $\ell=1,\dots,\ell_{\rm max}$, the effective sample size is (at least) $B\cdot S$. Formally, we aim to estimate $\bm{\theta}$ and test the hypotheses
$H_{0}:\langle\hat{\bm{A}}_{{f_{b}},s}\rangle=\bm{A}_{f_{b}}(\bm{\theta})\quad\hbox{vs.}$. $H_{a}:\langle\hat{\bm{A}}_{{f_{b}},s}\rangle\neq\bm{A}_{f_{b}}(\bm{\theta})$ (8)
@@ III.1 Parameter and covariance estimation
where $\Sigma_{{f_{b}},s}^{-1/2}$ is the inverse square-root matrix of $\Sigma_{{f_{b}},s}$ and can be obtained via standard procedures [36, 37]. The transformation in Eq. (10) is known as whitening or sphering and enables the construction of a statistically orthonormalized version of the angular power spectrum $\hat{\bm{A}}_{{f_{b}},s}$. In other words, under $H_{0}$ in Eq. (LABEL:eqn:test), $\tilde{\bm{A}}_{{f_{b}},s}$ has mean given by $\tilde{\bm{A}}_{{f_{b}},s}(\bm{\theta})$, its components are uncorrelated, and each has a unit variance.
@@
The solution of Eq. (11), can be shown to be consistent under $H_{0}$. That is, if the model $\bm{A}_{f_{b}}(\bm{\theta})$ is correctly specified, and if it is identifiable – that is, $\bm{A}_{f_{b}}(\bm{\theta})=\bm{A}_{f_{b}}(\bm{\theta}^{\prime})$ implies $\bm{\theta}=\bm{\theta}^{\prime}$ [Cf. 39, Ch. 7] – then, $\hat{\bm{\theta}}$ converges in probability [for a definition see 40, Ch. 2] to $\bm{\theta}_{0}$. Furthermore, under additional classical regularity conditions [Cf. 39, Ch. 7], $\hat{\bm{\theta}}$ is approximately Gaussian distributed with mean $\bm{\theta}_{0}$ and its variance approaches zero as $N\rightarrow\infty$. This can be seen by taking the derivative of the objective function in Eq. (11) with respect to $\bm{\theta}$, which shows that the generalized least squares estimator $\hat{\bm{\theta}}$ is a solution to the system of $p$ estimating equations:
@@
where $\dot{\bm{A}}_{{f_{b}},s}(\hat{\bm{\theta}})$ denotes the $\ell_{\rm max}\times p$ matrix with columns corresponding to the partial derivatives of $\tilde{\bm{A}}_{{f_{b}},s}(\bm{\theta})$ evaluated at $\hat{\bm{\theta}}$, i.e.,
$\dot{\bm{A}}_{{f_{b}},s}(\hat{\bm{\theta}})=\Sigma_{{f_{b}},s}^{-1/2}\left[\frac{\partial}{\partial\theta_{1}}\bm{A}_{f_{b}}(\bm{\theta}),\ldots,\frac{\partial}{\partial\theta_{p}}\bm{A}_{f_{b}}(\bm{\theta})\right]\biggl{|}_{\bm{\theta}=\hat{\bm{\theta}}},$
@@
When $\Sigma^{-1}_{{f_{b}},s}$ is unknown but a uniformly (for all $s$) consistent estimator is available, the statistical properties of $\hat{\bm{\theta}}$ described above remain valid provided that this estimator is substituted for $\Sigma^{-1}_{{f_{b}},s}$ in Eq. (10) [Cf. 39, Ch. 7].

### v3

@@ I Introduction
Estimating and testing models for the angular power spectrum is required in many astrophysical and cosmological settings, including the cosmic microwave background [2], galaxy sky surveys [3], and weak gravitational lensing surveys [4]. The focus here is on the stochastic gravitational-wave background (SGWB) resulting from the superposition of gravitational-wave (GW) signals that are too weak or too numerous to be resolved individually. While the SGWB and its anisotropies have not yet been detected by second-generation GW detectors such as the Advanced LIGO [5] and Advanced Virgo [6], upper limits have been placed on the GW energy density of the background as well as its angular power spectra [7, 8, 9, 10].
@@
As shown in Appendix A, the estimators of the angular power spectrum follow a generalized $\chi^{2}$ distribution, which cannot be accurately approximated by the Gaussian distribution. Moreover, using the generalized $\chi^{2}$ likelihood is nontrivial because, in general, the corresponding probability density function does not have a closed-form expression. Alternatively, one could devise a scheme that does not require any assumptions regarding the angular power spectrum’s underlying distribution. The latter approach is used here to implement and adapt the statistical procedure outlined in our companion paper [1] to the context of the SGWB.
@@ II Spherical Harmonics Search Methods for the SGWB
Data from ground-based interferometric gravitational-wave detectors may be used to generate estimators of the SGWB angular power spectrum. Each GW detector, labeled by $I=1,2$, produces a strain time series given by
$s_{I}(t)=h_{I}(t)+n_{I}(t),$ (1)
[…]
$h_{I}(t)=\int^{\infty}_{-\infty}\mathrm{d}f\int_{S^{2}}\mathrm{d}\hat{\Omega}h_{A}(f,\hat{\Omega})F_{I}^{A}(\hat{\Omega},t)e^{i2\pi f(t-\hat{\Omega}\cdot\vec{x}_{I}(t)/c)}.$ (2)
@@
[…]
$\langle h_{A}^{*}(f,\hat{\Omega})\,h_{A^{\prime}}(f^{\prime},\hat{\Omega}^{\prime})\rangle=\frac{1}{4}\,\mathcal{P}(f,\hat{\Omega})\,\delta_{AA^{\prime}}\,\delta(f-f^{\prime})\,\delta^{2}(\hat{\Omega},\hat{\Omega}^{\prime}),\,$ (4)
[…]
@@
$H(f)=\left(\frac{f}{f_{\rm ref}}\right)^{\alpha-3}.$ (6)
The factorization in Eq. (5) is performed under the assumption that the GW power is constant over each frequency band we consider, and it is used in other SGWB analyses, including broadband searches with bandwidths over 1700 Hz [25, 10]. We set the reference frequency $f_{\rm ref}=25$ Hz as is used in recent SGWB analyses [7, 10], and the spectral index $\alpha$ may be assigned a value depending on the type of SGWB model being considered. For the astrophysical SGWB resulting from merger events between compact objects such as black holes and neutron stars, $\alpha=2/3$ [7].
[…]
Note that these definitions assume that each GW frequency bin has the same anisotropy. This assumption can be relaxed, and the analysis can be pursued in separate frequency bins (or bands) $f_{b}$}. In this case, the spherical harmonic coefficients will become frequency dependent and we denote them as $a_{{f_{b}},\ell m}$ [34, 29].
The value of $\ell_{\rm max}$ determines the highest-order spherical harmonic modes that are considered in the analysis. This value may be chosen differently for different spectral models of the SGWB, and while $\ell_{\rm max}=4$ is typically used when $\alpha=2/3$ [7], $\ell_{\rm max}=8$ is used here. This choice of $\ell_{\rm max}$ is more appropriate for the frequency range over which we will evaluate the angular power spectrum. It has been shown that while $\ell_{\rm max}=4$ is an optimal choice for making a GW detection of the chosen SGWB model, it corresponds to an angular resolution that is lower than what is resolvable by the Advanced LIGO-Virgo detectors [35], which justifies considering spherical harmonic modes corresponding to a higher angular resolution.
@@
$\tilde{s}_{I}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t;f})\equiv\int^{t+\tau/2}_{t-\tau/2}dt^{\prime}e^{-i2\pi ft^{\prime}}s_{I}(t^{\prime}).$ (8)
@@
$\hat{C}(t;f)=\frac{2}{\tau}\tilde{s}_{I}^{*}(t;f)\tilde{s}_{J}(t;f),$ (9)
where, to simplify the notation, the $I,J$ indices corresponding to the two detectors from the cross-spectra terms have been dropped. Assuming that the noise of the two detectors is uncorrelated [25] [^1] While some sources of correlated noise such as Schumann resonances do exist, they do not affect searches for the SGWB given the current sensitivities of Advanced LIGO and Advanced Virgo [8, 36, 37]., the expectation value of the cross-correlation spectrum can be expressed as
$\langle\hat{C}(t;f)\rangle=\sum_{\ell=0}^{\ell_{\rm max}}\sum_{m=-\ell}^{\ell}H(f)\gamma_{\ell m}(t;f){a}_{{f_{b}},\ell m},$ (10)
@@
$N_{ft,f^{\prime}t^{\prime}}=\langle\hat{C}(t;f)\hat{C}^{*}(t^{\prime};f^{\prime})\rangle-\langle\hat{C}(t;f)\rangle\langle\hat{C}^{*}(t^{\prime};f^{\prime})\rangle$
@@
Here, $P_{I}(t;f)$ and $P_{J}(t;f)$ correspond to the one-sided power spectral densities of the two GW detectors’ outputs. Assuming the present GW signal powers (both cross- and auto-correlated) are much smaller than the detector noise power, we obtain the approximation in Eq. (11). For a given set of $a_{{f_{b}},\ell m}$, the likelihood function for the cross-correlation spectra in the frequency band $f_{b}$ is proportional to
$\exp\Bigl(-\sum_{{f},t}[\hat{C}(t;{f})-H({f})\,\sum_{\ell m}\gamma_{\ell m}(t;{f})\,{a}_{{f_{b}},\ell m}]^{*}$ (12)
$\quad N^{-1}_{{f}t,{f}t}[\hat{C}(t;{f})-H({f})\,\sum_{\ell m}\gamma_{\ell m}(t;{f})\,{a}_{{f_{b}},\ell m}]\Bigl)$
@@
Eq. (14) defines a map of the gravitational-wave background convolved with the antenna pattern of the detector pair, the so-called dirty map. The covariance matrix of the dirty map is given in Eq. (15) and is referred to as the Fisher matrix. Since Eq. (13) represents the deconvolution of the GW signal from the detector response, it is the so-called clean map. The covariance matrix of the clean map is obtained by inverting the covariance matrix of the dirty map. Detector insensitivity to certain sky directions manifests in small eigenvalues in the Fisher matrix that complicate its inversion. Such regularization is typically implemented via singular value decomposition to mitigate small singular values arising from directions in the sky to which the detector network is insensitive [25, 7]. However, this procedure introduces a bias that propagates the estimators and covariance of the angular power spectrum in non-trivial ways [29]. To avoid this bias, the Fisher matrix is not regularized in this study. Consequently, a lack of regularization will give increased uncertainties in estimators of the clean map and angular power spectrum [10].
Note that if Eq. (12) holds, each $\hat{a}_{{f_{b}},\ell m}$ is a linear combination of complex Gaussian variables; thus, they also follow a complex Gaussian distribution with mean $\langle\hat{a}_{{f_{b}},\ell m}\rangle={a}_{{f_{b}},\ell m}$ (discussed further below) and covariance $\operatorname{Cov}(\hat{a}_{{f_{b}},\ell m},\hat{a}_{{f_{b}},\ell^{\prime}m^{\prime}})=(\Gamma^{-1}_{{f_{b}}})_{\ell m,\ell^{\prime}m^{\prime}}.$ Indeed the same is true, at least approximately, even if Eq. (12) does not hold, but the summation over $t$ in Eq. (14) is taken over a large number of segments (as is the case with the Advanced LIGO and Advanced Virgo observing runs). This is true due to the central limit theorem as well as the assumption that the SGWB signal is weak relative to detector noise [24, 30, 38].
@@
From a statistical perspective, maximum likelihood is fundamentally a method for estimating deterministic quantities. This means the $\hat{a}_{{f_{b}},\ell m}$ in Eq. (13) serve as maximum likelihood estimators of the spherical harmonic coefficients, $a_{{f_{b}},\ell m}$, that uniquely characterize the SGWB of the only realization of the universe available. Therefore, similarly to what is implicitly done in many statistical analyses, the $a_{{f_{b}},\ell m}$ are not treated as random variables [e.g., 25, 39, 40, 7]. This is a reasonable assumption in practical scenarios, and treating the $a_{{f_{b}},\ell m}$ as fixed avoids the risk of incorrectly specifying their underlying distribution, which can substantially bias the resulting inference.
The proposed statistical methodology requires that multiple replicates of $\hat{a}_{{f_{b}},\ell m}$ are available. To ensure that is the case, we partition the total observation time $T$ into $S$ segments, each of duration $\Delta T$, so that $T=\Delta T\cdot S.$ We also divide the frequency range into $B$ discrete frequency bins $f_{b}$.
@@
One could always replace ${\Gamma}_{{f_{b}},s}^{-1}$ in Eq. (16) and Eq. (17) with its regularized counterpart if a regularization scheme was applied to invert the Fisher matrix [Cf. 25, 28, 41]. The derivation of Eq. (17) is given in Appendix B. Previous works [e.g., 28] have omitted the second term in the numerator of Eq. (17), implicitly making the assumption that $a_{{f_{b}},\ell m}$’s are complex Gaussian random variables with zero mean. However, this assumption is not consistent with estimating the coefficients $a_{{f_{b}},\ell m}$ (as realized in our universe) via the maximum likelihood technique – that is, likelihood maximization implicitly assumes the coefficients to be the (deterministic) limit to which the maximum likelihood estimates converge. In this manuscript, this inconsistency is avoided by treating $a_{{f_{b}},\ell m}$ as fixed.
@@ III Statistical Methods
$H_{0}:\langle\hat{\bm{A}}_{{f_{b}},s}\rangle=\bm{A}_{f_{b}}(\bm{\theta})\quad\hbox{vs.}$ $H_{a}:\langle\hat{\bm{A}}_{{f_{b}},s}\rangle\neq\bm{A}_{f_{b}}(\bm{\theta})$ (18)
@@ III.1 Parameter and covariance estimation
[…]
@@
The solution of Eq. (21), can be shown to be consistent [^4] Intuitively, consistency means that as more data are collected, the estimator $\widehat{\bm{\theta}}$ eventually becomes arbitrarily close to the true value $\theta_{0}$. A simple sufficient condition for consistency is that $\widehat{\theta}$ is asymptotically unbiased and its variance converges to zero. under $H_{0}$. That is, if the model $\bm{A}_{f_{b}}(\bm{\theta})$ is correctly specified, and if it is identifiable – i.e., $\bm{A}_{f_{b}}(\bm{\theta})=\bm{A}_{f_{b}}(\bm{\theta}^{\prime})$ implies $\bm{\theta}=\bm{\theta}^{\prime}$ [Cf. 45, Ch. 7] – then, $\hat{\bm{\theta}}$ converges in probability [for a definition see 46, Ch. 2] to $\bm{\theta}_{0}$. Furthermore, under additional classical regularity conditions [Cf. 45, Ch. 7], $\hat{\bm{\theta}}$ is approximately Gaussian distributed with mean $\bm{\theta}_{0}$ and its variance approaches zero as $N\rightarrow\infty$. This can be seen by taking the derivative of the function being minimized in Eq. (21) with respect to $\bm{\theta}$, which shows that the generalized least squares estimator $\hat{\bm{\theta}}$ solves the system of $p$ estimating equations:
@@
[…]
$\dot{\bm{A}}_{{f_{b}},s}(\hat{\bm{\theta}})=\Sigma_{{f_{b}},s}^{-1/2}\left[\frac{\partial}{\partial\theta_{1}}\bm{A}_{f_{b}}(\bm{\theta}),\ldots,\frac{\partial}{\partial\theta_{p}}\bm{A}_{f_{b}}(\bm{\theta})\right]\biggl|_{\bm{\theta}=\hat{\bm{\theta}}}$, (23)
@@
When $\Sigma^{-1}_{{f_{b}},s}$ is unknown but a uniformly (for all $s$) consistent estimator is available, the statistical properties of $\hat{\bm{\theta}}$ described above remain valid provided that this estimator is substituted for $\Sigma^{-1}_{{f_{b}},s}$ in Eq. (20) [Cf. 45, Ch. 7].